Exercise 19-8 Express Delivery is a rapidly growing delivery service. Last year, 80% of its revenue came from the delivery of mailing pouches and small, standardized delivery boxes (which provides a 20% contribution margin). The other 20% of its revenue came from delivering non-standardized boxes (which provides a 70% contribution margin), with the rapid growth of Internet retail sales, Express believes that there are great opportunities for growth in the delivery of non-standardized boxes. The company has fixed costs of $12,987,000. (a) What is the companys break-even point in total sales dollars? At the break-even point, how much of the companys sales are provided by each type of service? (Use Weighted-Average Contribution Margin Ratio rounded to 4 decimal places e.g. 0.2552 and round final answers to 0 decimal places, e.g. 2,510.) Total break-even sales Sale of mail pouches and small boxes Sale of non-standard boxes (b) The companys management would like to hold its fixed costs constant but shift its sales mix so that 60% of its revenue comes from the delivery of non-standardized boxes and the remainder from pouches and small boxes. If this were to occur, what would be the companys break even sales, and what amount of sales would be provided by each service type? (Use Weighted-Average Contribution Margin Ratio rounded to 4 decimal places e.g. 0.2552 and round final answers to 0 decimal places, e.g. 2,510.) Total break-even sales Sale of mail pouches and small boxes Sale of non-standardized boxes Click if you would like to Show Work for this question:

Please make all the answers clear in order (maybe bold them?). Thank you!!!
Answer #1
a
Weighted average contribution margin = (80%*20%)+(20%*70%)= 30%
Total break even sales=Fixed costs/CM ratio = 12987000/30%= 43290000
Sale of mail pouches and small boxes = 43290000*80%=34632000
Sale of non standard boxes = 43290000*20%=8658000
b
Weighted average contribution margin = (40%*20%)+(60%*70%)= 50%
Total break even sales=Fixed costs/CM ratio = 12987000/50%= 25974000
Sale of mail pouches and small boxes = 25974000*40%=10389600
Sale of non standard boxes = 25974000*60%=15584400